OpenStudy (anonymous):

6(4z-1)-5(z+3)=8(z+1)

OpenStudy (anonymous):

you start by distributing, 24z - 6 - 5z - 15 = 8z + 8 combining like terms we get, 19z - 21= 8z + 8 combining more like terms we get, 11z = 29 divide by 11 z = 29/11

OpenStudy (anonymous):